I was absent the other day and I need help on the Pythagorean throrem. My teacher didn't explain it to me. My question is how do begin to find the square root for a problem like a=5 b= ? c= 7 and c is the hypotenuse. Please help.
Answer by flame8855(424)

7^2 = 5^2 + b^2
b^2= 49-25 = 24
b = sqrt(24)